The toxic gas, hydrogen fluoride, HF, is produced from the double displacement reaction of calcium fluoride, CaF2, and concentrated sulfuric acid H2SO4. Starting with 10.0 g of CaF2 and 15.5 g of H2SO4, how many grams of HF will be produced? Which is limiting reactant? What grams of the excess reactant will be left after the reaction?
1
Expert's answer
2020-10-26T14:48:26-0400

CaF2 + H2SO4 → 2HF + CaSO4

M(CaF2) = 78.07 g/mol

M(H2SO4) = 98.08 g/mol

M(HF) = 20.01 g/mol

n = m/M

n(CaF2) "= \frac{10.0}{78.07} = 0.128 \;mol"

n(H2SO4)"= \frac{15.5}{98.08} = 0.158 \;mol"

Proportion according to the reaction:

1 mol CaF2 : 1 mol H2SO4

Actual amount of reactants:

0.128 mol CaF2 : 0.158 mol H2SO4

So, H2SO4 is in excess and CaF2 is the limiting reactant.

n(HF) = 2n(CaF2) "= 2 \times 0.128 = 0.256 \;mol"

m(CaF2) "= 0.256 \times 20.01 = 5.12 \;g"

n(H2SO4) = 0.128 mol (needs for reaction)

∆n(H2SO4) = 0.158 – 0.128 = 0.03 mol

Mass of excess reactant:

m(H2SO4) "= 0.03 \times 98.08 = 2.94 \;g"